how do you solve 13x-49=-179 how do i solve for x
Answer by nerdybill(7384)

13x-49=-179
.
idea is to "isolate" the unknown to one side of the equation:
13x-49=-179
start by adding 49 to both sides:
13x-49+49 = -179+49
13x = -130
divide both sides by 13:
13x/13 = -130/13
x = -10